It takes you 12 minutes to go 2.5 miles, what is your speed in miles per hour?
Answer by ReadingBoosters(3246)

Convert minutes to hours
12 min x 1 hr/60 min = 1/5 hr
2.5 mi / (1/5 hr) = mi/hr
2.5 / .2 = 12.5 mi/hr